OpenStudy (izzy18970):

I need help solving this 2!!!! 13=g/3 + 4 -1/2x - 7= -11

undeadknight26 (undeadknight26):

im thinking: 13 = g/3 + 4 -4 -4 8 = g/3 x3 x3 g =24?

undeadknight26 (undeadknight26):

the second one would be -.5x - 7 = -11 +7 +7 -.5x = -4 x.-5 x-.5 x = 2?

undeadknight26 (undeadknight26):

I believe this is it...

OpenStudy (izzy18970):

when you check the answer to see if it is right, I don't get 13=13,/

OpenStudy (izzy18970):

how do you check it though

undeadknight26 (undeadknight26):

lol i have no idea.. @BangkokGarrett

OpenStudy (izzy18970):

very funny

OpenStudy (izzy18970):

oh well, thx for the help